# Multipoint distributions of the KPZ fixed point with compactly supported initial conditions

### v2

@@ 3.2 Multipoint distribution of TASEP with general initial configurations
The following theorem is essentially [Liu22, Theorem 2.1], where the integrand $\mathcal{D}_{Y}(z_{1},\cdots,z_{m-1})$ is defined by a Fredholm determinant formula or a series expansion formula and the initial condition information is encoded in a characteristic function of $Y$. It was proved that the choice of characteristic functions does not affect the value of $\mathcal{D}_{Y}$ function. We will present the formula with a new characteristic function $\mathrm{ch}_{Y}$ as in (3.5) in Theorem 3.4 that is suitable for asymptotic analysis.
@@
Given $Y=(\cdots<y_{2}<y_{1})\in\mathbb{Z}^{\mathbb{Z}_{+}}$. Consider TASEP with initial particle locations $X_{0}=Y$. Let $(k_{1},t_{1}),\cdots,(k_{m},t_{m})$ be $m$ distinct points in $\mathbb{Z}_{+}\times\mathbb{R}_{+}$. Then for any integers $a_{1},\cdots,a_{m}$,
$\mathbb{P}_{Y}\left(\bigcap_{\ell=1}^{m}\{\mathrm{x}_{k_{\ell}}(t_{\ell})\geq a_{\ell}\}\right)=\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{1}}{2\pi\mathrm{i}z_{1}(1-z_{1})}\cdots\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{m-1}}{2\pi\mathrm{i}z_{m-1}(1-z_{m-1})}\mathcal{D}_{Y}(z_{1},\cdots,z_{m-1}),$ (3.11)
where $\mathbb{P}_{Y}$ denotes the probability given $X(0)=Y$. The function $\mathcal{D}_{Y}(z_{1},\cdots,z_{m-1})$ is defined as a Fredholm determinant in Definition 3.6 , or equivalently as a series in Definition 3.7.
3.2.1 Fredholm determinant representation of $\mathcal{D}_{Y}(z_{1},\cdots,z_{m-1})$
The definition of $\mathcal{D}_{Y}(z_{1},\cdots,z_{m-1})$ is very similar to its limiting counterpart $\mathbf{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1})$ defined in Section 2.1 and 2.2, either as a Fredholm determinant $\det(\mathrm{I}-\mathcal{K}_{1}\mathcal{K}_{Y})$ or as a Fredholm series expansion. We will only use the series expansion formula in this paper but we present both formulas here for completeness and possible later uses.
@@ 3.2.1.1 Spaces of the operators
We will define the operators on two specific spaces of nested contours with complex measures depending on $\bm{z}=(z_{1},\cdots,z_{m-1})$, where $z_{\ell}\neq 1$ for each $1\leq\ell\leq m-1$. Recall the definition of the two regions $\Omega_{\mathrm{L}}$ and $\Omega_{\mathrm{R}}$ from (3.1).
Suppose $\Sigma_{m,\mathrm{L}}^{\mathrm{out}},\cdots,\Sigma_{2,\mathrm{L}}^{\mathrm{out}}$, $\Sigma_{1,\mathrm{L}}$, $\Sigma_{2,\mathrm{L}}^{\mathrm{in}},\cdots,\Sigma_{m,\mathrm{L}}^{\mathrm{in}}$ are 2m12m-1 nested simple closed contours, from outside to inside, in $\Omega_{\mathrm{L}}$ enclosing the point 1-1. Similarly, $\Sigma_{m,\mathrm{R}}^{\mathrm{out}},\cdots,\Sigma_{2,\mathrm{R}}^{\mathrm{out}}$, $\Sigma_{1,\mathrm{R}}$, $\Sigma_{2,\mathrm{R}}^{\mathrm{in}},\cdots,\Sigma_{m,\mathrm{R}}^{\mathrm{in}}$ are 2m12m-1 nested simple closed contours, from outside to inside, in $\Omega_{\mathrm{R}}$ enclosing the point 0. See Figure 2 for an illustration of the contours. These contours are all counterclockwise oriented.
@@
$\Sigma_{\ell,\mathrm{L}}:=\Sigma_{\ell,\mathrm{L}}^{\mathrm{out}}\cup\Sigma_{\ell,\mathrm{L}}^{\mathrm{in}},\qquad\Sigma_{\ell,\mathrm{R}}:=\Sigma_{\ell,\mathrm{R}}^{\mathrm{out}}\cup\Sigma_{\ell,\mathrm{R}}^{\mathrm{in}},\qquad\ell=2,\cdots,m,$ (3.12)
@@
${\mathrm{d}}\mu(w)={\mathrm{d}}\mu_{\bm{z}}(w):=\begin{dcases}\frac{-z_{\ell-1}}{1-z_{\ell-1}}\frac{{\mathrm{d}}w}{2\pi{\mathrm{i}}},&w\in\Sigma_{\ell,\mathrm{L}}^{\mathrm{out}}\cup\Sigma_{\ell,\mathrm{R}}^{\mathrm{out}},\quad\ell=2,\cdots,m,\\ \frac{1}{1-z_{\ell-1}}\frac{{\mathrm{d}}w}{2\pi{\mathrm{i}}},&w\in\Sigma_{\ell,\mathrm{L}}^{\mathrm{in}}\cup\Sigma_{\ell,\mathrm{R}}^{\mathrm{in}},\quad\ell=2,\cdots,m,\\ \frac{{\mathrm{d}}w}{2\pi{\mathrm{i}}},&w\in\Sigma_{1,\mathrm{L}}\cup\Sigma_{1,\mathrm{R}}.\end{dcases}$
@@ 3.2.1.2 Operators $\mathcal{K}_{1}$ and $\mathcal{K}_{Y}$
Now we introduce the operators $\mathcal{K}_{1}$ and $\mathcal{K}_{Y}$ to define $\mathcal{D}_{Y}(z_{1},\cdots,z_{m-1})$ in Theorem 3.5. We assume that $\bm{z}=(z_{1},\cdots,z_{m-1})$ is the same as in Section 3.2.1.1. Let
@@
$\mathcal{D}_{Y}(z_{1},\cdots,z_{m-1})=\det\left(\mathrm{I}-\mathcal{K}_{1}\mathcal{K}_{Y}\right),$
@@
$f_{i}(w):=\begin{dcases}w^{k_{i}-k_{i-1}}(w+1)^{-(a_{i}-a_{i-1})-(k_{i}-k_{i-1})}\mathrm{e}^{(t_{i}-t_{i-1})w},&i=2,\cdots,m,\\ w^{k_{1}}(w+1)^{-a_{1}-k_{1}}\mathrm{e}^{t_{1}w},&i=1,\end{dcases}$ (3.15)
@@
3.2.2 Series expansion formula for $\mathcal{D}_{Y}(z_{1},\cdots,z_{m-1})$
@@
$\mathcal{D}_{Y}(z_{1},\cdots,z_{m-1}):=\sum_{\bm{n}\in(\mathbb{Z}_{\geq 0})^{m}}\frac{1}{(\bm{n}!)^{2}}\mathcal{D}_{Y}^{(\mathbf{n})}(z_{1},\cdots,z_{m-1}),$ (3.16)
with $\bm{n}!=n_{1}!\cdots n_{m}!$ for $\bm{n}=(n_{1},\cdots,n_{m})$. Here
$\begin{split}&\mathcal{D}^{(\mathbf{n})}_{Y}(z_{1},\cdots,z_{m-1})=\mathcal{D}^{(\mathbf{n})}_{Y}(z_{1},\cdots,z_{m-1};(x_{1},t_{1},a_{1}),\cdots,(x_{m},t_{m},a_{m}))\\ &=\prod_{\ell=1}^{m-1}\left(1-z_{\ell}\right)^{n_{\ell}}\left(1-z_{\ell}^{-1}\right)^{n_{\ell+1}}\left(\prod_{\ell=1}^{m}\prod_{i_{\ell}=1}^{n_{\ell}}\int_{\Sigma_{\ell,\mathrm{L}}}\mathrm{d}\mu_{\bm{z}}(u_{i_{\ell}}^{(\ell)})\int_{\Sigma_{\ell,\mathrm{R}}}\mathrm{d}\mu_{\bm{z}}(v_{i_{\ell}}^{(\ell)})\right)\prod_{\ell=1}^{m}\prod_{i_{\ell}=1}^{n_{\ell}}\frac{f_{\ell}(u_{i_{\ell}}^{(\ell)})}{f_{\ell}(v_{i_{\ell}}^{(\ell)})}\\ &\quad\cdot\det\begin{bmatrix}\mathrm{ch}_{Y}(v^{(1)}_{i},u^{(1)}_{j})\end{bmatrix}_{1\leq i,j\leq n_{1}}\cdot\prod_{\ell=1}^{m-1}\mathrm{C}\left(U^{(\ell)}\sqcup V^{(\ell+1)};V^{(\ell)}\sqcup U^{(\ell+1)}\right)\cdot\mathrm{C}(U^{(m)};V^{(m)}),\end{split}$ (3.17)
with $U^{(\ell)}=(u_{1}^{(\ell)},\cdots,u_{n_{\ell}}^{(\ell)})$ and $V^{(\ell)}=(v_{1}^{(\ell)},\cdots,v_{n_{\ell}}^{(\ell)})$, and the functions $f_{\ell}$ defined in (3.15) for $1\leq\ell\leq m$.
@@ 4 Convergence of the TASEP formula
In this section we will take proper scaling limit of the TASEP formulas (see Theorem 3.5), to get the corresponding KPZ fixed point formulas. We start with the setup for the proper rescaling.
@@ 4.1 TASEP height function and $1:2:3$ rescaling
The TASEP particle configurations can be encoded into the corresponding height functions $H(x,t)$ defined as the unique function $\mathbb{R}\times\mathbb{R}_{+}\to\mathbb{R}$ satisfying the following conditions:
@@
The dynamics of the height function is as follows. Each local maximum of the height function turns into a local minimum after an independent exponential time with rate $1$. After each flip of max to min the height at the flip descreases by $2$ while the height at the other integer points remain unchanged. The values at general $x\in\mathbb{R}$ are then determined by linear interpolations. Note that here we follow the convention in [MQR21] where the height function decreases in time, instead of increasing as in some other literature. More explicitly, for the TASEP Markov chain $X_{t}=(\mathrm{x}_{i}(t))_{i\geq 1}$ we define its height function as
@@
In particular the initial height function $\mathsf{h}$ corresponding to the initial particle configuration $Y$ is
@@
Now we introduce the proper rescaling for the TASEP height function so that it will converge to the KPZ fixed point. For $\varepsilon>0$ we define the rescaled TASEP height function $\mathcal{H}^{\varepsilon}(\alpha,\tau)$ for $(\alpha,\tau)\in\mathbb{R}\times\mathbb{R}_{+}$ as follows:
@@
In particular $\mathcal{H}^{\varepsilon}(\alpha,0)=\varepsilon^{\frac{1}{2}}\cdot H(2\varepsilon^{-1}\alpha,0)=:\mathfrak{h}^{\varepsilon}(\alpha)$. It was shown in [MQR21, Theorem 3.13] that if $\mathfrak{h}^{\varepsilon}\to\mathfrak{h}$ in $\mathrm{UC}$ as $\varepsilon\to 0$, then for any positive integer $m$ one has $(\mathcal{H}^{\varepsilon}(\cdot,\tau_{1};\mathfrak{h}^{\varepsilon}),\cdots,\mathcal{H}^{\varepsilon}(\cdot,\tau_{m};\mathfrak{h}^{\varepsilon}))$ converges in distribution to $(\mathcal{H}(\cdot,\tau_{1};\mathfrak{h}),\cdots,\mathcal{H}(\cdot,\tau_{m};\mathfrak{h}))$ in the topology of $\mathrm{UC}^{m}$, where $\mathcal{H}(\cdot,\cdot;\mathfrak{h})$ is the KPZ fixed point starting from the initial condition $\mathfrak{h}$. This in particular implies

### v3

@@ 3.2 Multipoint distribution of TASEP with general initial configurations
As we discussed in the previous subsection, the multipoint distribution of TASEP only depends on a finite number of rightmost particles. Hence, in this subsection, we consider the TASEP model with finitely many particles.
The following theorem is essentially [Liu22, Theorem 2.1], where the integrand $\mathcal{D}_{Y}(z_{1},\ldots,z_{m-1})$ is defined by a Fredholm determinant formula or a series expansion formula and the initial condition information is encoded in a characteristic function of $Y$. It was proved that the choice of characteristic functions does not affect the value of the $\mathcal{D}_{Y}$ function. We will present the formula with a new characteristic function $\mathrm{ch}_{Y}$ as in (3.5) in Theorem 3.4 that is suitable for asymptotic analysis.
@@
Given $Y=(y_{N},\ldots,y_{2},y_{1})\in\mathbb{Z}^{N}$ satisfying $y_{N}<\cdots<y_{1}$, where $N$ is a positive integer. Consider TASEP with initial particle locations $X_{0}=Y$. Let $(k_{1},t_{1}),\ldots,(k_{m},t_{m})$ be $m$ distinct points in $\{1,\ldots,N\}\times\mathbb{R}_{+}$ satisfying $t_{1}\leq t_{2}\leq\cdots\leq t_{m}$. Then for any integers $a_{1},\ldots,a_{m}$,
$\mathbb{P}_{Y}\left(\bigcap_{\ell=1}^{m}\{\mathrm{x}_{k_{\ell}}(t_{\ell})\geq a_{\ell}\}\right)=\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{1}}{2\pi\mathrm{i}z_{1}(1-z_{1})}\cdots\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{m-1}}{2\pi\mathrm{i}z_{m-1}(1-z_{m-1})}\mathcal{D}_{Y}(z_{1},\ldots,z_{m-1}),$ (3.12)
where $\mathbb{P}_{Y}$ denotes the probability given $X(0)=Y$. The function $\mathcal{D}_{Y}(z_{1},\ldots,z_{m-1})$ is defined as a Fredholm determinant in Definition 3.6, or equivalently as a series in Definition 3.7.
3.2.1 Fredholm determinant representation of $\mathcal{D}_{Y}(z_{1},\ldots,z_{m-1})$
The definition of $\mathcal{D}_{Y}(z_{1},\ldots,z_{m-1})$ is very similar to its limiting counterpart $\mathbf{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1})$ defined in Section 2.1 and 2.2, either as a Fredholm determinant $\det(\mathrm{I}-\mathcal{K}_{1}\mathcal{K}_{Y})$ or as a Fredholm series expansion. We will only use the series expansion formula in this paper, but we present both formulas here for completeness and possible later uses.
@@ 3.2.1.1 Spaces of the operators
We will define the operators on two specific spaces of nested contours with complex measures depending on $\bm{z}=(z_{1},\ldots,z_{m-1})$, where $z_{\ell}\neq 1$ for each $1\leq\ell\leq m-1$. Recall the definition of the two regions $\Omega_{\mathrm{L}}$ and $\Omega_{\mathrm{R}}$ from (3.1).
Suppose $\Sigma_{m,\mathrm{L}}^{\mathrm{out}},\ldots,\Sigma_{2,\mathrm{L}}^{\mathrm{out}}$, $\Sigma_{1,\mathrm{L}}$, $\Sigma_{2,\mathrm{L}}^{\mathrm{in}},\ldots,\Sigma_{m,\mathrm{L}}^{\mathrm{in}}$ are 2m12m-1 nested simple closed contours, from outside to inside, in $\Omega_{\mathrm{L}}$ enclosing the point 1-1. Similarly, $\Sigma_{m,\mathrm{R}}^{\mathrm{out}},\ldots,\Sigma_{2,\mathrm{R}}^{\mathrm{out}}$, $\Sigma_{1,\mathrm{R}}$, $\Sigma_{2,\mathrm{R}}^{\mathrm{in}},\ldots,\Sigma_{m,\mathrm{R}}^{\mathrm{in}}$ are 2m12m-1 nested simple closed contours, from outside to inside, in $\Omega_{\mathrm{R}}$ enclosing the point 0. See Figure 2 for an illustration of the contours. These contours are all counterclockwise oriented.
@@
$\Sigma_{\ell,\mathrm{L}}:=\Sigma_{\ell,\mathrm{L}}^{\mathrm{out}}\cup\Sigma_{\ell,\mathrm{L}}^{\mathrm{in}},\qquad\Sigma_{\ell,\mathrm{R}}:=\Sigma_{\ell,\mathrm{R}}^{\mathrm{out}}\cup\Sigma_{\ell,\mathrm{R}}^{\mathrm{in}},\qquad\ell=2,\ldots,m,$ (3.13)
@@
${\mathrm{d}}\mu(w)={\mathrm{d}}\mu_{\bm{z}}(w):=\begin{dcases}\frac{-z_{\ell-1}}{1-z_{\ell-1}}\frac{{\mathrm{d}}w}{2\pi{\mathrm{i}}},&w\in\Sigma_{\ell,\mathrm{L}}^{\mathrm{out}}\cup\Sigma_{\ell,\mathrm{R}}^{\mathrm{out}},\quad\ell=2,\ldots,m,\\ \frac{1}{1-z_{\ell-1}}\frac{{\mathrm{d}}w}{2\pi{\mathrm{i}}},&w\in\Sigma_{\ell,\mathrm{L}}^{\mathrm{in}}\cup\Sigma_{\ell,\mathrm{R}}^{\mathrm{in}},\quad\ell=2,\ldots,m,\\ \frac{{\mathrm{d}}w}{2\pi{\mathrm{i}}},&w\in\Sigma_{1,\mathrm{L}}\cup\Sigma_{1,\mathrm{R}}.\end{dcases}$
@@ 3.2.1.2 Operators $\mathcal{K}_{1}$ and $\mathcal{K}_{Y}$
Now we introduce the operators $\mathcal{K}_{1}$ and $\mathcal{K}_{Y}$ to define $\mathcal{D}_{Y}(z_{1},\ldots,z_{m-1})$ in Theorem 3.5. We assume that $\bm{z}=(z_{1},\ldots,z_{m-1})$ is the same as in Section 3.2.1.1. Let
@@
$\mathcal{D}_{Y}(z_{1},\ldots,z_{m-1})=\det\left(\mathrm{I}-\mathcal{K}_{1}\mathcal{K}_{Y}\right),$
@@
$f_{i}(w):=\begin{dcases}w^{k_{i}-k_{i-1}}(w+1)^{-(a_{i}-a_{i-1})-(k_{i}-k_{i-1})}\mathrm{e}^{(t_{i}-t_{i-1})w},&i=2,\ldots,m,\\ w^{k_{1}}(w+1)^{-a_{1}-k_{1}}\mathrm{e}^{t_{1}w},&i=1,\end{dcases}$ (3.16)
@@
3.2.2 Series expansion formula for $\mathcal{D}_{Y}(z_{1},\ldots,z_{m-1})$
@@
$\mathcal{D}_{Y}(z_{1},\ldots,z_{m-1}):=\sum_{\bm{n}\in(\mathbb{Z}_{\geq 0})^{m}}\frac{1}{(\bm{n}!)^{2}}\mathcal{D}_{Y}^{(\mathbf{n})}(z_{1},\ldots,z_{m-1}),$ (3.17)
with $\bm{n}!=n_{1}!\cdots n_{m}!$ for $\bm{n}=(n_{1},\ldots,n_{m})$. Here
$\begin{split}&\mathcal{D}^{(\mathbf{n})}_{Y}(z_{1},\ldots,z_{m-1})=\mathcal{D}^{(\mathbf{n})}_{Y}(z_{1},\ldots,z_{m-1};(x_{1},t_{1},a_{1}),\ldots,(x_{m},t_{m},a_{m}))\\ &=\prod_{\ell=1}^{m-1}\left(1-z_{\ell}\right)^{n_{\ell}}\left(1-z_{\ell}^{-1}\right)^{n_{\ell+1}}\left(\prod_{\ell=1}^{m}\prod_{i_{\ell}=1}^{n_{\ell}}\int_{\Sigma_{\ell,\mathrm{L}}}\mathrm{d}\mu_{\bm{z}}(u_{i_{\ell}}^{(\ell)})\int_{\Sigma_{\ell,\mathrm{R}}}\mathrm{d}\mu_{\bm{z}}(v_{i_{\ell}}^{(\ell)})\right)\prod_{\ell=1}^{m}\prod_{i_{\ell}=1}^{n_{\ell}}\frac{f_{\ell}(u_{i_{\ell}}^{(\ell)})}{f_{\ell}(v_{i_{\ell}}^{(\ell)})}\\ &\quad\cdot\det\begin{bmatrix}\mathrm{ch}_{Y}(v^{(1)}_{i},u^{(1)}_{j})\end{bmatrix}_{1\leq i,j\leq n_{1}}\cdot\prod_{\ell=1}^{m-1}\mathrm{C}\left(U^{(\ell)}\sqcup V^{(\ell+1)};V^{(\ell)}\sqcup U^{(\ell+1)}\right)\cdot\mathrm{C}(U^{(m)};V^{(m)}),\end{split}$ (3.18)
with $U^{(\ell)}=(u_{1}^{(\ell)},\ldots,u_{n_{\ell}}^{(\ell)})$ and $V^{(\ell)}=(v_{1}^{(\ell)},\ldots,v_{n_{\ell}}^{(\ell)})$, and the functions $f_{\ell}$ defined in (3.16) for $1\leq\ell\leq m$.
@@ 4 Convergence of the TASEP formula
In this section, we will take the proper scaling limit of the TASEP formulas (see Theorem 3.5), to get the corresponding KPZ fixed point formulas. We start with the setup for the proper rescaling.
@@ 4.1 TASEP height function and $1:2:3$ rescaling
The TASEP particle configuration can be encoded into the corresponding height function $H(x,t)$, defined as the unique function $\mathbb{R}\times\mathbb{R}_{+}\to\mathbb{R}$ satisfying the following conditions:
@@
The dynamics of the height function is as follows. Each local maximum of the height function turns into a local minimum after an independent exponential time with rate $1$. After each flip of max to min, the height at the flip decreases by $2$ while the height at the other integer points remain unchanged. The values at general $x\in\mathbb{R}$ are then determined by linear interpolations. Note that here we follow the convention in [MQR21] where the height function decreases in time, instead of increasing as in some other literature. More explicitly, for the TASEP Markov chain $X_{t}=(\mathrm{x}_{i}(t))_{i\geq 1}$ we define its height function as
@@
In particular, the initial height function $\mathsf{h}$ corresponding to the initial particle configuration $Y$ is
@@
Now we introduce the proper rescaling for the TASEP height function so that it will converge to the KPZ fixed point. For $\varepsilon>0$, we define the rescaled TASEP height function $\mathcal{H}^{\varepsilon}(\alpha,\tau)$ for $(\alpha,\tau)\in\mathbb{R}\times\mathbb{R}_{+}$ as follows:
@@
In particular $\mathcal{H}^{\varepsilon}(\alpha,0)=\varepsilon^{\frac{1}{2}}\cdot H(2\varepsilon^{-1}\alpha,0)=:\mathfrak{h}^{\varepsilon}(\alpha)$. It was shown in [MQR21, Theorem 3.13] that if $\mathfrak{h}^{\varepsilon}\to\mathfrak{h}$ in $\mathrm{UC}$ as $\varepsilon\to 0$, then for any positive integer $m$ one has $(\mathcal{H}^{\varepsilon}(\cdot,\tau_{1};\mathfrak{h}^{\varepsilon}),\ldots,\mathcal{H}^{\varepsilon}(\cdot,\tau_{m};\mathfrak{h}^{\varepsilon}))$ converges in distribution to $(\mathcal{H}(\cdot,\tau_{1};\mathfrak{h}),\ldots,\mathcal{H}(\cdot,\tau_{m};\mathfrak{h}))$ in the topology of $\mathrm{UC}^{m}$, where $\mathcal{H}(\cdot,\cdot;\mathfrak{h})$ is the KPZ fixed point starting from the initial condition $\mathfrak{h}$. This in particular implies